# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v2

@@ Appendix E SkipSponge Threshold Hyperparameter Study
Figure 12. Accuracy for ResNet18 during the SkipSponge attack for different thresholds.
Appendix F Fired Percentage Graphs SkipSponge
@@
Figure 13. We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and VGG16 after the SkipSponge attack. The average is calculated over all images in the validation set.
Figure 14. We display the differences in the average percentage of positive activations in all layers for a clean ResNet18 model and ResNet18 after the SkipSponge attack. The average is calculated over all images in the validation set.
Appendix G Fired Percentage Graphs Sponge Poisoning
In Figure 15, we show the average fired percentage on the validation set per layer for VGG16 and for ResNet18 in Figure 16. Graphs are for Sponge Poisoning with $\lambda=2.5$, $p=0.05$ and $\sigma=1e-04$. In the figures, we see that sponge causes a fired percentage increase in all layers of VGG16 and ResNet18 for all datasets, typically by 10% or more. We hypothesize that this is because Sponge Poisoning benefits from including the classification loss during training and can maintain the performance of the model by adjusting other parameters in the model to compensate for the increased bias values in the batch normalization layers.
Figure 15. We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and a Sponge Poisoned VGG16 model. The average is calculated over all images in the validation set.
Figure 16. The differences in the average percentage of positive activations in all layers for a clean ResNet18 model and a Sponge Poisoned ResNet18 model. The average is calculated over all images in the validation set.

### v3

@@ Appendix A SkipSponge Threshold Hyperparameter Study
Figure 13: Accuracy for ResNet18 during the SkipSponge attack for different thresholds.
A-A Fired Percentage Graphs SkipSponge
@@
Figure 14: We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and VGG16 after the SkipSponge attack. The average is calculated over all images in the validation set.
Figure 15: We display the differences in the average percentage of positive activations in all layers for a clean ResNet18 model and ResNet18 after the SkipSponge attack. The average is calculated over all images in the validation set.
A-B Fired Percentage Graphs Sponge Poisoning
In Figure 16, we show the average fired percentage on the validation set per layer for VGG16 and for ResNet18 in Figure 17. Graphs are for Sponge Poisoning with $\lambda=2.5$, $\delta=0.05$ and $\sigma=1e-04$. In the figures, we see that sponge causes a fired percentage increase in all layers of VGG16 and ResNet18 for all datasets, typically by 10% or more. We hypothesize that this is because Sponge Poisoning benefits from including the classification loss during training and can maintain the performance of the model by adjusting other parameters in the model to compensate for the increased bias values in the batch normalization layers.
Figure 16: We display the differences in the average percentage of positive activations in all layers for a clean VGG16 model and a Sponge Poisoned VGG16 model. The average is calculated over all images in the validation set.
Figure 17: The differences in the average percentage of positive activations in all layers for a clean ResNet18 model and a Sponge Poisoned ResNet18 model. The average is calculated over all images in the validation set.